# Beyond Winning Strategies: Admissible and Admissible Winning Strategies for Quantitative Reachability Games

### v2

@@ B.3 Proof of Lemma 2
On simplifying, we get $\operatorname{aVal}(h,\sigma^{\prime})\geq\operatorname{aVal}(h)>\operatorname{cVal}(h,\sigma)$. This statement implies that there exists $\tau\in\mathrm{T}$ such that $\operatorname{Val}(h\cdot P^{h}(\sigma^{\prime},\tau))>\operatorname{Val}(h\cdot P^{h}(\sigma,\tau))$. Note that since, $\operatorname{Val}(h\cdot P^{h}(\sigma,\tau))=\operatorname{Val}(h)+\operatorname{Val}(P^{v_{s}}(\sigma,\tau))$ we get $\operatorname{Val}(P^{v_{s}}(\sigma^{\prime},\tau))>\operatorname{Val}(P^{v_{s}}(\sigma,\tau))$. This contradicts the assumption that $\sigma^{\prime}$ dominates $\sigma$ as $\sigma^{\prime}$ should always have a payoff that is equal to or lower than $\sigma$.
@@ B.4 Proof of Theorem 1
For the play induced by $\sigma_{1}$, $\operatorname{cVal}(v_{0},\sigma_{1})<\operatorname{aVal}(v_{0})$, $\operatorname{cVal}(v_{0}v_{1}v_{3},\sigma_{1})<\operatorname{aVal}(v_{0}v_{1}v_{3}),\ldots\operatorname{cVal}(v_{0}v_{1}v_{3}v_{6}v_{7},\sigma_{1})<\operatorname{aVal}(v_{0}v_{1}v_{3}v_{6}v_{7})$. Thus, $\operatorname{cVal}(h,\sigma_{1})<\operatorname{aVal}(h)$ for all the histories $h$ compatible with $\sigma_{1}$. For the play induced by $\sigma_{2}$, $\operatorname{cVal}(v_{0},\sigma_{2})<\operatorname{aVal}(v_{0})$, but $\operatorname{cVal}(v_{0}v_{1}v_{3},\sigma_{2})>\operatorname{aVal}(v_{0}v_{1}v_{3})$. Notice that $\operatorname{cVal}(v_{0}v_{1}v_{3}v_{6},\sigma_{2})<\operatorname{aVal}(v_{0}v_{1}v_{3}v_{6})$ and thus we need to check admissibility for all histories $h$ compatible with $\sigma$. Hence, for history $h:=v_{0}v_{1}v_{3}$, $\sigma_{2}$ is not $\operatorname{\mathbf{SC}}$. If $h=v_{0}v_{2}v_{4}v_{6}$, then strategy $\sigma_{3}$ and $\sigma_{4}$ are both admissible. Thus, $\sigma_{3}$ and $\sigma_{4}$ are both $\operatorname{\mathbf{SC}}$ strategies. This proves that $\operatorname{\mathbf{SC}}$ strategies are history-dependent. ∎
@@ B.5 Proof of Lemma 3
$\sigma^{\prime}$ is not $\operatorname{\mathbf{WCO}}$: This implies that $\operatorname{aVal}(h)\neq\operatorname{aVal}(h,\sigma^{\prime})$. As $\sigma^{\prime}$ is not $\operatorname{\mathbf{WCO}}$, it implies that there exists an adversarial strategy $\tau\in\mathrm{T}$ for which the payoff $\operatorname{Val}(h\cdot P^{v_{s}}(\sigma^{\prime},\tau))>\operatorname{Val}(h\cdot P^{v_{s}}(\sigma,\tau))$. Since, $\operatorname{Val}(h\cdot P^{v_{s}}(\sigma,\tau))=\operatorname{Val}(h)+\operatorname{Val}(P^{v_{s}}(\sigma,\tau))$, we get $\operatorname{Val}(P^{v_{s}}(\sigma^{\prime},\tau))>\operatorname{Val}(P^{v_{s}}(\sigma,\tau))$. This contradicts our assumption that $\sigma^{\prime}\succ\sigma$ as $\sigma^{\prime}$ should never do worse than $\sigma$.
@@ 5       $W_{win}^{\prime}=W_{win}$
7             $W_{win}(v)=\max(C(v,a)+W^{\prime}_{win}(v’))\quad\text{if}\;v\in V_{e}$
@@
8             $W_{win}(v)=\min(C(v,a)+W^{\prime}_{win}(v’))\quad\text{if}\;v\in V_{e}$
9             $W_{win}(v)=\min(C(v,a)+W^{\prime}_{win}(v’))\quad\text{if}\;v\in V_{s}$
10             $\sigma_{win}(v)=\operatorname*{arg\,min}_{a}(C(v,a)+W^{\prime}_{win}(v’))\quad\text{if}\;v\in V_{s}$
@@ B.6 Proof of Theorem 2
In the adversarial setting, Khachiyan et al. [2008] show that memoryless strategies are sufficient for two-player, non-negative weights scenario. Thus, we have that $\operatorname{cVal}(h,\sigma)=\operatorname{cVal}(\operatorname{last}(h),\sigma)$ for $\operatorname{\mathbf{Co-Op}}$ and $\operatorname{aVal}(h,\sigma)=\operatorname{aVal}(\operatorname{last}(h),\sigma)$ for $\operatorname{\mathbf{WCO}}$.
Next, we show that for a given history $h$, memoryless strategies are sufficient to be optimal adversarial-cooperative. We show this by proving that $\operatorname{cVal}(\operatorname{last}(h),\sigma)=\operatorname{acVal}(h)$. Given history $h$, we define $\operatorname{aVal}(\mathcal{G},h)$ and $\operatorname{cVal}(\mathcal{G},h)$ to be the optional adversarial and cooperative value in $\mathcal{G}$. We extend the definition to $\operatorname{acVal}(\mathcal{G},h)$ accordingly. We denote by $\Sigma(\mathcal{G})$ and $\mathrm{T}(\mathcal{G})$ the set of all valid strategies in $\mathcal{G}$.
Given $\mathcal{G}$, history $h$, we define $\mathcal{G}^{\prime}$ to be the subgame such that the initial state in $\mathcal{G}^{\prime}$ is $\operatorname{last}(h)$ and every state in $\mathcal{G}^{\prime}$ satisfies condition $\operatorname{aVal}(\mathcal{G},v)\leq\operatorname{aVal}(\mathcal{G},h)\;\forall v\in V$. We note that $\operatorname{aVal}(\mathcal{G},v_{f})=\operatorname{cVal}(\mathcal{G},v_{f})=0$ and $\operatorname{aVal}(\mathcal{G},v)\geq\operatorname{cVal}(\mathcal{G},v)\neq 0\;\forall v\notin V_{f}$. Thus, every subgame $\mathcal{G}^{\prime}$ will include the goal states $v\in V_{f}$. As $\mathcal{G}^{\prime}$ is the subgame of $\mathcal{G}$ the strategies in $\mathcal{G}^{\prime}$ can be uniquely mapped to $\mathcal{G}$. Notice that $\Sigma(\mathcal{G}^{\prime})\subseteq\Sigma(\mathcal{G})$ where $\Sigma(\mathcal{G}^{\prime})$ is the set of all valid strategies in $\mathcal{G}^{\prime}$. The weights of the edges in $\mathcal{G}^{\prime}$ are the same as $\mathcal{G}$.
@@ B.9 Proof of Theorem 4
The result follows from Thm. 3 that states that $\operatorname{\mathbf{SC}}$ or $\operatorname{\mathbf{WCo-Op}}$ are necessary and sufficient conditions for a strategy to be admissible. Using the results of [Brihaye et al., 2017, Thm. 1], we show that the value iteration algorithm on $\mathcal{G}$ will converge after finite iterations to a fixed point. Then, using [Brihaye et al., 2017, Corollary 18 and Proposition 19], we prove that witnessing strategies for $\operatorname{\mathbf{WCO}}$ and $\operatorname{\mathbf{Co-Op}}$ always exists.
Further, from Thm. 2, we know that for any history $h$, $\operatorname{acVal}(h)$ is same as the optimal cooperative value in the sub-game starting from $\operatorname{last}(h)$. Thus, witnessing strategies for $\operatorname{\mathbf{WCO}}$ and $\operatorname{\mathbf{Co-Op}}$ are sufficient conditions for $\operatorname{\mathbf{WCo-Op}}$ always exist.
@@ B.12 Proof of Theorem 5
Let us define $\operatorname{aVal}(\mathcal{G},v)$ and $\operatorname{cVal}(\mathcal{G},v)$ to be the positional adversarial and cooperative value for state $v$ in $\mathcal{G}$. We extend the definition to $\operatorname{acVal}(\mathcal{G},v)$ accordingly. We begin our proof by first noting that $\mathcal{G}^{\prime}$ is a finite tree arena where the leaf nodes are partitioned into goal states for the Sys player and sink states. By construction, every state that is not a leaf node in $\mathcal{G}^{\prime}$ has at least one outgoing edge. Only plays that reach a goal state in $\mathcal{G}^{\prime}$ have a finite payoff value. Plays that fail to reach a goal state have a payoff of +$\infty$+ ∞. Thus, using Khachiyan et al. [2008]’s algorithm, we can compute optimal $\operatorname{aVal}(\mathcal{G}^{\prime},v)$ and $\operatorname{cVal}(\mathcal{G}^{\prime},v)$ and the corresponding witnessing strategies. Thus, game $\mathcal{G}^{\prime}$ is well-formed, which implies $\operatorname{\mathbf{SC}}$ and $\operatorname{\mathbf{WCo-Op}}$ always exists (from Thm. 4). Hence, there always exists an admissible strategy in $\mathcal{G}^{\prime}$.
@@ B.17 Proof of Theorem 8
Alg. 2 is the Value Iteration algorithm for computing optimal $\operatorname{Val}$s and optimal strategy from every state in $\mathcal{G}$ Brihaye et al. [2017][Adapted and Modified from Algo. 1]. Line 2 and Line 2 correspond to computation when the Env player is playing adversarially (maximally) and cooperatively (minimally), respectively. When Env is playing adversarially, Alg. 2 returns optimal $\operatorname{aVal}$ and worst-case optimal strategy $\operatorname{\mathbf{WCO}}$. For $\operatorname{aVal}$ computation, $\sigma_{win}$ is the optimal winning strategy from every state in $V_{win}$.
@@ Appendix C Note on $\operatorname{\mathbf{SCO}}$ vs $\operatorname{\mathbf{SC}}$
Brenguier et al. [2016] define strongly-cooperative optimal strategies in order to show that every $\operatorname{\mathbf{SCO}}$ is admissible. Then, they proceed to shows that $\operatorname{\mathbf{SCO}}$ strategies always exists in determined (well-formed as per their terminology) games, i.e., you can always find a strategy $\sigma$ such that $\operatorname{aVal}(h)=\operatorname{aVal}(h,\sigma)$ and $\operatorname{cVal}(h)=\operatorname{cVal}(h,\sigma)$ for every $h\in\operatorname{Plays}^{v}(\sigma)$.

### v3

@@ 6 Conclusion
Acknowledgments
This work was supported in part by the University of Colorado Boulder and by Strategic University Research Partnership (SURP) grants from the NASA Jet Propulsion Laboratory (JPL) (RSA 1688009 and 1704147). The authors thank the reviewers for their constructive feedback.
@@ B.3 Proof of Lemma 2
On simplifying, we get $\operatorname{aVal}(h,\sigma^{\prime})\geq\operatorname{aVal}(h)>\operatorname{cVal}(h,\sigma)$. This statement implies that there exists $\tau\in\mathrm{T}$ such that $\operatorname{Val}(h\cdot P^{h}(\sigma^{\prime},\tau))>\operatorname{Val}(h\cdot P^{h}(\sigma,\tau))$. Note that since, $\operatorname{Val}(h\cdot P^{h}(\sigma,\tau))=\operatorname{Val}(h^{\prime})+\operatorname{Val}(P^{v_{s}}(\sigma,\tau))$ where $h^{\prime}=v_{0}\ldots v_{|h|-2}$. Thus, we get $\operatorname{Val}(P^{v_{s}}(\sigma^{\prime},\tau))>\operatorname{Val}(P^{v_{s}}(\sigma,\tau))$. This contradicts the assumption that $\sigma^{\prime}$ dominates $\sigma$ as $\sigma^{\prime}$ should always have a payoff that is equal to or lower than $\sigma$.
@@ B.4 Proof of Theorem 1
For the play induced by $\sigma_{1}$, $\operatorname{cVal}(v_{0},\sigma_{1})<\operatorname{aVal}(v_{0})$, $\operatorname{cVal}(v_{0}v_{1}v_{3},\sigma_{1})<\operatorname{aVal}(v_{0}v_{1}v_{3})$, $\operatorname{cVal}(v_{0}v_{1}v_{3}v_{6}v_{7},\sigma_{1})<\operatorname{aVal}(v_{0}v_{1}v_{3}v_{6}v_{7})$. Thus, $\operatorname{cVal}(h,\sigma_{1})<\operatorname{aVal}(h)$ for all the histories $h$ compatible with $\sigma_{1}$. For the play induced by $\sigma_{2}$, $\operatorname{cVal}(v_{0},\sigma_{2})<\operatorname{aVal}(v_{0})$, but $\operatorname{cVal}(v_{0}v_{1}v_{3},\sigma_{2})>\operatorname{aVal}(v_{0}v_{1}v_{3})$. Notice that $\operatorname{cVal}(v_{0}v_{1}v_{3}v_{6},\sigma_{2})<\operatorname{aVal}(v_{0}v_{1}v_{3}v_{6})$ and thus we need to check admissibility for all histories $h$ compatible with $\sigma$. Hence, for history $h:=v_{0}v_{1}v_{3}$, $\sigma_{2}$ is not $\operatorname{\mathbf{SC}}$. If $h=v_{0}v_{2}v_{4}v_{6}$, then strategy $\sigma_{3}$ and $\sigma_{4}$ are both admissible. Thus, $\sigma_{3}$ and $\sigma_{4}$ are both $\operatorname{\mathbf{SC}}$ strategies. This proves that $\operatorname{\mathbf{SC}}$ strategies are history-dependent. ∎
@@ B.5 Proof of Lemma 3
$\sigma^{\prime}$ is not $\operatorname{\mathbf{WCO}}$: This implies that $\operatorname{aVal}(h)\neq\operatorname{aVal}(h,\sigma^{\prime})$. As $\sigma^{\prime}$ is not $\operatorname{\mathbf{WCO}}$, it implies that there exists an adversarial strategy $\tau\in\mathrm{T}$ for which the payoff $\operatorname{Val}(h\cdot P^{v_{s}}(\sigma^{\prime},\tau))>\operatorname{Val}(h\cdot P^{v_{s}}(\sigma,\tau))$. Since, $\operatorname{Val}(h\cdot P^{v_{s}}(\sigma,\tau))=\operatorname{Val}(h^{\prime})+\operatorname{Val}(P^{v_{s}}(\sigma,\tau))$, where $h^{\prime}=v_{0}\dots v_{|h|-2}$. Thus, we get $\operatorname{Val}(P^{v_{s}}(\sigma^{\prime},\tau))>\operatorname{Val}(P^{v_{s}}(\sigma,\tau))$. This contradicts our assumption that $\sigma^{\prime}\succ\sigma$ as $\sigma^{\prime}$ should never do worse than $\sigma$.
@@ 5       $W_{win}^{\prime}=W_{win}$
7             $W_{win}(v)=\max(C(v,a)+W^{\prime}_{win}(v^{\prime}))\quad\text{if}\;v\in V_{e}$
@@
8             $W_{win}(v)=\min(C(v,a)+W^{\prime}_{win}(v^{\prime}))\quad\text{if}\;v\in V_{e}$
9             $W_{win}(v)=\min(C(v,a)+W^{\prime}_{win}(v^{\prime}))\quad\text{if}\;v\in V_{s}$
10             $\sigma_{win}(v)=\operatorname*{arg\,min}_{a}(C(v,a)+W^{\prime}_{win}(v^{\prime}))\quad\text{if}\;v\in V_{s}$
@@ B.6 Proof of Theorem 2
In the adversarial setting, Khachiyan et al. show that memoryless strategies are sufficient for two-player, non-negative weights scenario.
Next, we show that for a given history $h$, memoryless strategies are sufficient to be optimal adversarial-cooperative. We show this by proving that $\operatorname{cVal}(\operatorname{last}(h),\sigma)=\operatorname{acVal}(h)$. Given history $h$, we define $\operatorname{aVal}(\mathcal{G},h)$ and $\operatorname{cVal}(\mathcal{G},h)$ to be the optimal adversarial and cooperative value in $\mathcal{G}$. We extend the definition to $\operatorname{acVal}(\mathcal{G},h)$ accordingly. We denote by $\Sigma(\mathcal{G})$ and $\mathrm{T}(\mathcal{G})$ the set of all valid strategies in $\mathcal{G}$.
Given $\mathcal{G}$ and history $h$, we define $\mathcal{G}^{\prime}$ to be the subgame such that the initial state in $\mathcal{G}^{\prime}$ is $\operatorname{last}(h)$ and every state in $\mathcal{G}^{\prime}$ satisfies condition $\operatorname{aVal}(\mathcal{G},v)\leq\operatorname{aVal}(\mathcal{G},h)\;\forall v\in V$. We note that $\operatorname{aVal}(\mathcal{G},v_{f})=\operatorname{cVal}(\mathcal{G},v_{f})=0$ and $\operatorname{aVal}(\mathcal{G},v)\geq\operatorname{cVal}(\mathcal{G},v)\neq 0\;\forall v\notin V_{f}$. Thus, every subgame $\mathcal{G}^{\prime}$ will include the goal states $v\in V_{f}$. As $\mathcal{G}^{\prime}$ is the subgame of $\mathcal{G}$ the strategies in $\mathcal{G}^{\prime}$ can be uniquely mapped to $\mathcal{G}$. Notice that $\Sigma(\mathcal{G}^{\prime})\subseteq\Sigma(\mathcal{G})$ where $\Sigma(\mathcal{G}^{\prime})$ is the set of all valid strategies in $\mathcal{G}^{\prime}$. The weights of the edges in $\mathcal{G}^{\prime}$ are the same as $\mathcal{G}$.
@@ B.9 Proof of Theorem 4
The result follows from Thm. 3 that states that $\operatorname{\mathbf{SC}}$ or $\operatorname{\mathbf{WCo-Op}}$ are necessary and sufficient conditions for a strategy to be admissible. Using the results of Brihaye et al., Thm. 1, we show that the value iteration algorithm on $\mathcal{G}$ will converge after finite iterations to a fixed point. Then, using Brihaye et al., Corollary 18 and Proposition 19, we prove that witnessing strategies for $\operatorname{\mathbf{WCO}}$ and $\operatorname{\mathbf{Co-Op}}$ always exists.
Further, from Thm. 2, we know that for any history $h$, $\operatorname{acVal}(h)$ is same as the optimal cooperative value in the sub-game starting from $\operatorname{last}(h)$. Thus, witnessing strategies for $\operatorname{\mathbf{WCO}}$ and $\operatorname{\mathbf{Co-Op}}$ are sufficient conditions for $\operatorname{\mathbf{WCo-Op}}$ to always exist.
@@ B.12 Proof of Theorem 5
Let us define $\operatorname{aVal}(\mathcal{G},v)$ and $\operatorname{cVal}(\mathcal{G},v)$ to be the positional adversarial and cooperative value for state $v$ in $\mathcal{G}$. We extend the definition to $\operatorname{acVal}(\mathcal{G},v)$ accordingly. We begin our proof by first noting that $\mathcal{G}^{\prime}$ is a finite tree arena where the leaf nodes are partitioned into goal states for the Sys player and sink states. By construction, every state that is not a leaf node in $\mathcal{G}^{\prime}$ has at least one outgoing edge. Only plays that reach a goal state in $\mathcal{G}^{\prime}$ have a finite payoff value. Plays that fail to reach a goal state have a payoff of +$\infty$+ ∞. Thus, using Khachiyan et al.’s algorithm, we can compute optimal $\operatorname{aVal}(\mathcal{G}^{\prime},v)$ and $\operatorname{cVal}(\mathcal{G}^{\prime},v)$ and the corresponding witnessing strategies. Thus, game $\mathcal{G}^{\prime}$ is well-formed, which implies $\operatorname{\mathbf{SC}}$ and $\operatorname{\mathbf{WCo-Op}}$ always exists (from Thm. 4). Hence, there always exists an admissible strategy in $\mathcal{G}^{\prime}$.
@@ B.17 Proof of Theorem 8
Alg. 2 is the Value Iteration algorithm for computing optimal $\operatorname{Val}$s and optimal strategy from every state in $\mathcal{G}$ Brihaye et al. [2017] (Adapted and Modified from Alg. 1). Line 2 and Line 2 correspond to computation when the Env player is playing adversarially (maximally) and cooperatively (minimally), respectively. When Env is playing adversarially, Alg. 2 returns optimal $\operatorname{aVal}$ and worst-case optimal strategy $\operatorname{\mathbf{WCO}}$. For $\operatorname{aVal}$ computation, $\sigma_{win}$ is the optimal winning strategy from every state in $V_{win}$.
@@ Appendix C Note on $\operatorname{\mathbf{SCO}}$ vs $\operatorname{\mathbf{SC}}$
Brenguier et al. define strongly-cooperative optimal strategies in order to show that every $\operatorname{\mathbf{SCO}}$ is admissible. Then, they proceed to shows that $\operatorname{\mathbf{SCO}}$ strategies always exists in determined (well-formed as per their terminology) games, i.e., you can always find a strategy $\sigma$ such that $\operatorname{aVal}(h)=\operatorname{aVal}(h,\sigma)$ and $\operatorname{cVal}(h)=\operatorname{cVal}(h,\sigma)$ for every $h\in\operatorname{Plays}^{v}(\sigma)$.